# PRIME: Scaffolding Manipulation Tasks with Behavior Primitives for Data-Efficient Imitation Learning

### v2

@@ III-B2 Dynamic Programming Algorithm
Specifically, considering a task demonstration of the target task denoted as $\tau=(s_{0},a_{0},...,s_{\mathcal{T}})$ where $a_{t}$ represents low-level motor actions. We define the objective function in our dynamic programming process as $f(i)$, which represents the probability of decomposing $(s_{0},a_{0},...,s_{i})$ into an optimal sequence of primitives $\left(s_{0},(p_{0},x_{0}),s_{j_{1}},(p_{j_{1}},x_{j_{1}}),...,s_{i}\right)$. We iteratively update the objective function
$f(i)=\max\limits_{t<i,p,x}f(t)\cdot(\alpha\cdot\text{IDM}(p,x|s_{t},s_{i}))$ (1)
by maximizing the product of probabilities of primitive sequences. We multiply a factor $\alpha$ to $\text{IDM}(p,x|s_{t},s_{i})$, where $\alpha=0.0001$ if $p=\texttt{other}$; otherwise, $\alpha=1$. We use the factor $\alpha$ to discourage the trivial mapping of motor actions to atomic primitive instances. This encourages the trajectory parser to parse demonstrations into primitive sequences composed solely of non-atomic primitives, typically characterized by shorter sequences. Upon completing the dynamic programming, we can extract the optimal primitive sequences denoted as $(s_{0},p_{0},x_{0},s_{t_{1}},p_{1},x_{1},\ldots,s_{t_{\mathcal{M}}}=s_{\mathcal{T}})$ from the final value $f(\mathcal{T})$, where $\mathcal{M}$ represents the length of the segmented sequence.
@@ III-C Policy Learning
Given that the segmented primitive sequences are notably shorter than the task demonstrations, leading to significantly smaller segmented data than the size of the demonstrations, we introduce a stepwise augmentation technique to enrich the segmented data and increase the scope of supervision for imitation learning. We assume that beginning from any point within the demonstration, the decomposition will remain consistent in subsequent segments. In other words, we presume that the start and end points of each segment in the parsed sequences are the same across all “suffixes” of demonstrations, where a suffix is defined as a portion of the demonstration beginning from any intermediate state and continuing to the final state. With this assumption, for each segmented primitive sequences $\tau^{\prime}=(s_{0},(p_{0},x_{0}),s_{t_{1}},(p_{1},x_{1}),...,s_{\mathcal{T}})$ parsed from a task demonstration $\tau=(s_{0},a_{0},...,s_{\mathcal{T}})$, we can get an augmented tuple $(s_{l},p^{\prime}_{d},x^{\prime}_{d},s_{t_{d+1}})$ at each timestep $l$, where $s_{t_{d}}<l<s_{t_{d+1}}$, $(p^{\prime}_{d},x^{\prime}_{d})=\operatorname*{argmax}\text{IDM}(.|s_{l},s_{t_{d+1}})$. We incorporate these augmented tuples $\{(s_{l},p^{\prime}_{d},x^{\prime}_{d})\}$ into the training dataset of policy.
@@ III-D Implementation Details
We use the primitives from Nasiriany et al. [31] to implement our library of primitives with minor modifications. We implement the following five primitives:
  • Reaching: The robot moves its end-effector to a target location $(x,y,z)$ and yaw angle $\psi$.
@@
  • Pushing: The robot reaches a starting location $(x,y,z)$ at a yaw angle $\psi$ and then moves its end-effector by a displacement $(\delta_{x},\delta_{y},\delta_{z})$.
  • Atomic: The robot applies a robot motor action.
To reduce the complexity of the mapping, we factorize the IDM into a primitive IDM $\text{IDM}_{\text{prim}}(p|s,s^{\prime})$ capable of identifying primitive types and a parameter IDM $\text{IDM}_{\text{param}}(x|s,s^{\prime},p)$ capable of predicting parameters of the primitive, where
$\text{IDM}(p,x|s,s^{\prime})=\text{IDM}_{\text{prim}}(p|s,s^{\prime})\cdot\text{IDM}_{\text{param}}(x|s,s^{\prime},p)$ (2)
Our primitive IDM and parameter IDM are two separate networks. Both networks take a pair of observations, $s$ and $s^{\prime}$, and encode these observations with a pair of ResNet-18 encoders. The primitive IDM outputs a Softmax distribution over primitives, while the parameter IDM outputs a Gaussian mixture model over parameter values. Similarly to the IDM, we factorize the policy $\pi(p,x|s)$ into a primitive policy $\pi_{\text{prim}}(p|s)$ and a parameter policy $\pi_{\text{param}}(x|s,p)$. During deployment, the primitive policy first predicts the next primitive type $p$ given the current state $s$. Following this, the parameter policy predicts the corresponding parameters $x$ given inputs $s$ and $p$. We train these two policy networks separately with behavioral cloning. These networks take a single observation as input (the current observation $s$) and have architectures and output spaces similar to their counterparts in the IDM. We use the robomimic framework [27] to implement and train these networks.
@@ IV Experiments
Our experiments are designed to answer the following questions: 1) How does erform for imitation learning in low-data regimes? 2) Which design choices are critical to 3) How effective is the trajectory parser in 4) Is easible for practical deployment to real-world robot tasks?
Figure 3: Simulated Tasks. We perform evaluations on three tasks from the robosuite simulator [56]. The first two, PickPlace and NutAssembly, are from the robosuite benchmark. We introduce a third task, TidyUp, to study long-horizon tasks and test the inverse dynamics model’s generalization to unseen environments. We create four environment variants in this domain, denoted as (A, B, C, D). TidyUp task is designed in environment (D), and we collect human demonstrations for TidyUp in the same environment (D). To gauge the inverse dynamics model’s generalization capability, we train two IDMs: IDM-D, based solely on data from environment (D), and IDM-ABC, trained on data from environments (A, B, C). While IDM-D is our default model for experiments, we use IDM-ABC to evaluate generalization in unseen environments.
@@ IV-A1 Manipulation Tasks
We validate our approach and examine the above questions in simulated and real-world tasks. We perform evaluations in three simulated tasks from the robosuite simulator [56] (see Fig. 3) and two real-world tasks (see Fig. 6):
@@
  • CleanUp. This domain replicates the real-world experimental setup introduced by Nasiriany et al. [31], involving two tasks: CleanUp-Bin and CleanUp-Stack. The tasks require the robot to push a popcorn box to a serving area and move butter to a target location. See Fig. 6 for the tasks.
We employ a Franka Emika Panda robot with operational space control for all tasks. In simulation and the real world, we opt for 5-DoF control: 3 for end-effector position $(\delta_{x},\delta_{y},\delta_{z})$, 1 for yaw orientation, and 1 for the gripper. The agent observes wrist-camera and third-person images and receives proprioceptive data. Vision-based policies are trained for all tasks, with only 30 human demonstrations per task. For the training data of IDM, we collect 1M transitions for PickPlace, 3M transitions for NutAssembly, and 5M transitions for TidyUp. Notably, all tasks are completed based on the same set of primitives.
@@ IV-A3 Baselines
We contrast our method with three imitation learning baselines: BC-RNN [27], its fine-tuning variant BC-RNN (FT), and SAILOR [30]. BC-RNN uses LSTM as the backbone network architecture and learns through behavioral cloning. BC-RNN (FT) pre-trains on prior data and fine-tunes with target task demonstrations. SAILOR pre-trains a skill encoder to extract skill latents from prior data and learns the policy using task demonstrations and pre-trained skill representations through imitation learning. The prior data for pre-training BC-RNN (FT) and SAILOR consists of all low-level transitions in the training dataset of IDM in Experimental Results
TABLE I: Success rates in TidyUp for ablation studies.
Ours No Pretraining Greedy Algorithm No Placing
$\mathbf{0.989\pm 0.005}$ $0.944\pm 0.003$ $0.859\pm 0.004$ $0.0\pm 0$ 0.0
@@
IV-A4 Quantitative Results
Fig. 4 demonstrates our method’s substantial superiority over all baselines, achieving success rates exceeding 95% across all tasks with remarkable robustness. This showcases the effectiveness of our approach in achieving data-efficient imitation learning through the decomposition of raw sensory demonstrations into concise primitive sequences. Furthermore, our policy often attempts the same primitive type and slightly adjusts primitive parameters after a failure. The results indicate that our primitive policy is more capable of learning this recovery attempt following a failure than baseline policies, which require predicting a sequence of actions for recovery.
IV-A5 Ablation Studies
We perform ablation studies in the long-horizon TidyUp task to study the importance of design choices in the overall performance of our framework. We compare our methods with ablations: 1) without policy pretraining (No Pretraining), 2) utilizing greedy algorithm to interpret demonstrations into primitive sequences instead of dynamic programming (Greedy Algorithm), where the greedy algorithm selects next primitive by examining all states in the demonstration that come after the current state and choosing the one with the highest probability, 3) removing the placing primitive from the primitive library (No Placing). As shown in Table I, omitting policy pre-training or substituting dynamic programming with the greedy algorithm leads to decreased performance, highlighting their essential roles. Additionally, when the primitive set lacks a placing primitive, task completion becomes unattainable. This emphasizes the need for a comprehensive set of primitives that encompass all necessary behaviors. Nasiriany et al. [31] also demonstrated a similar finding that removing primitives significantly hurts performance.
IV-A6 Model Analysis
@@
IV-B Real-World Evaluation
@@
We evaluate the performance of gainst an imitation learning baseline (BC-RNN) on two real-world CleanUp task variants: CleanUp-Bin and CleanUp-Stack. To ensure safe data collection, we execute our self-supervised data collection procedure in simulation for training an IDM and apply the learned IDM to interpreting real-world demonstrations. The observations in real-world demonstrations contain camera images, object pose information, and robot proprioceptive states. We build a state-based IDM that takes object pose information and robot proprioceptive states as its input and directly transfers it to parsing real-world demonstrations that also contain object pose information. We highlight that a single IDM is utilized for segmenting demonstrations across two tasks as both tasks fall within the same domain. Furthermore, leveraging the presence of camera images in demonstrations, we develop a visual-based policy for predicting motor actions based on camera images and robot states. To obtain object pose information on the real robot, we use a pose estimator [48].
The results in Table IV show our method’s notable advantage over BC-RNN, primarily due to its effective task scaffolding. Similar to results in the simulation, our method also exhibits robust recovery behavior on real robots. For instance, the robot persistently attempts to grasp the target object until it succeeds. Our method has a success rate below 70% in CleanUp-Stack, versus 90% in CleanUp-Stack, mainly because stacking demands greater precision and sim2real training for IDM leads to errors.
@@ V Conclusion
We presented a data-efficient imitation learning approach that decomposes task demonstrations into sequences of primitives and leverages imitation learning to acquire the high-level control policy for sequencing parameterized primitives. We leverage a self-supervised data collection procedure involving the random execution of primitives within the environment. We evaluate oth in the simulation and on the real robot, demonstrating significant improvements over baseline approaches in low-data regimes. While emonstrates significant results, a promising direction for future research lies in learning a scalable library of low-level primitives and composing these diverse primitives. This approach holds the potential to facilitate curriculum learning, allowing for the progressive acquisition of increasingly complex tasks. A limitation of this study is the use of sim2real experiments for IDM training, which may not suit challenging real-world tasks. Extending IDM training to real-world settings is left for future research.
@@ ACKNOWLEDGMENT
We thank Yifeng Zhu for his generous help with the real robot infrastructure. We also thank Jake Grigsby, Mingyo Seo, Rutav Shah, and Zhenyu Jiang for their valuable feedback on this manuscript. This work has been partially supported by the National Science Foundation (EFRI-2318065, FRR-2145283), the Office of Naval Research (N00014-22-1-2204), UT Good Systems, and the Machine Learning Laboratory.

### v3

@@ III-B2 Dynamic Programming Algorithm
Specifically, considering a task demonstration denoted as $\tau=(s_{0},a_{0},...,s_{\mathcal{T}})$ where $a_{t}$ represents low-level motor actions. We define the objective function in our dynamic programming process as $f(i)$, which represents the probability of decomposing $(s_{0},a_{0},...,s_{i})$ into an optimal sequence of primitives $\left(s_{0},(p_{0},x_{0}),s_{j_{1}},(p_{1},x_{1}),...,s_{i}\right)$. We iteratively update the objective function
$f(i)=\max\limits_{p,x,t<i}f(t)\cdot(\alpha\cdot\text{IDM}(p,x|s_{t},s_{i}))$ (1)
by maximizing the product of probabilities of primitive sequences. We multiply a factor $\alpha$ to $\text{IDM}(p,x|s_{t},s_{i})$, where $\alpha$ is a small constant ($\alpha=0.0001$ in our implementation) if $p$ is in “other” category; otherwise, $\alpha=1$. We use the factor $\alpha$ to penalize mappings to the category “other”. Upon completing the dynamic programming, we can extract the optimal primitive sequences denoted as $(s_{0},p_{0},x_{0},s_{t_{1}},p_{1},x_{1},\ldots,s_{t_{\mathcal{M}}}=s_{\mathcal{T}})$ from the final value $f(\mathcal{T})$, where $\mathcal{M}$ represents the length of the segmented sequence.
@@ III-C Policy Learning
Given that the segmented primitive sequences are notably shorter than the task demonstrations, leading to significantly smaller segmented data than the size of the demonstrations, we introduce a stepwise augmentation technique to enrich the segmented data and increase the scope of supervision for imitation learning. We assume that beginning from any point within the demonstration, the decomposition will remain consistent in subsequent segments. In other words, we presume that start and end points of each segment in parsed sequences are the same across all “suffixes” of demonstrations, where a suffix is defined as a portion of the demonstration beginning from any intermediate state and continuing to the final state. With this assumption, for each segmented primitive sequences $\tau^{\prime}=(s_{0},(p_{0},x_{0}),s_{t_{1}},(p_{1},x_{1}),...,s_{\mathcal{T}})$ parsed from a task demonstration $\tau=(s_{0},a_{0},...,s_{\mathcal{T}})$, we can get an augmented tuple $(s_{l},p^{\prime}_{d},x^{\prime}_{d},s_{t_{d+1}})$ at each timestep $l$, where $s_{t_{d}}<l<s_{t_{d+1}}$, $(p^{\prime}_{d},x^{\prime}_{d})=\operatorname*{argmax}\text{IDM}(.|s_{l},s_{t_{d+1}})$. We incorporate these augmented tuples $\{(s_{l},p^{\prime}_{d},x^{\prime}_{d})\}$ into the training dataset of policy.
@@ III-D Implementation Details
We use the primitives from Nasiriany et al. [nasiriany2022augmenting] to implement our library of primitives with minor modifications. These task-independent, hard-coded APIs can be directly adapted to new situations within the same domain, as these APIs only require robot proprioceptive information as input. We implement the following four primitives:
  • Reaching: The robot moves its end-effector to a target location $(x,y,z)$ and yaw angle $\psi$ in a collision-free path.
@@
  • Pushing: The robot reaches a starting location $(x,y,z)$ at a yaw angle $\psi$ in a collision-free path and then moves its end-effector by a displacement $(\delta_{x},\delta_{y},\delta_{z})$.
To reduce the complexity of the mapping, we factorize the IDM into a primitive IDM, i.e. $\text{IDM}_{\text{prim}}(p|s,s^{\prime})$, capable of identifying primitive types and a parameter IDM, i.e. $\text{IDM}_{\text{param}}(x|s,s^{\prime},p)$, capable of predicting parameters of the primitive, where
$\text{IDM}(p,x|s,s^{\prime})=\text{IDM}_{\text{prim}}(p|s,s^{\prime})\cdot\text{IDM}_{\text{param}}(x|s,s^{\prime},p).$ (2)
Our primitive IDM and parameter IDM are two separate networks. Both networks take a pair of observations, $s$ and $s^{\prime}$, and encode these observations with a pair of ResNet-18 encoders. A 2-layer MLP follows the image encoder in both networks. The primitive IDM outputs a Softmax distribution over primitives, while the parameter IDM outputs a Gaussian mixture model over parameter values. Our model architecture is similar to the default architecture in the RoboMimic framework [mandlekar2021matters], as the paper reported that ResNet encoder and GMM policy uniformly perform better than other design choices. Similarly to the IDM, we factorize the policy $\pi(p,x|s)$ into a primitive policy, i.e. $\pi_{\text{prim}}(p|s)$, and a parameter policy, i.e. $\pi_{\text{param}}(x|s,p)$. During deployment, the primitive policy first predicts the next primitive type $p$ given the current state $s$. Following this, the parameter policy predicts the corresponding parameters $x$ given inputs $s$ and $p$. We train these two policy networks separately with behavioral cloning. These networks take a single observation as input (the current observation $s$) and have architectures and output spaces similar to their counterparts in the IDM. We use RoboMimic [mandlekar2021matters] to implement and train these networks.
@@ IV Experiments
Our experiments are designed to answer the following questions: 1) How does PRIME perform for imitation learning in low-data regimes? 2) Which design choices are critical to PRIME? 3) How effective is the trajectory parser in PRIME? 4) Is PRIME feasible for practical deployment to real-world robot tasks? 5) Are learned skills compatible with PRIME?
Figure 3: Simulated Tasks. We perform evaluations on three tasks from the RoboSuite simulator [zhu2020robosuite]. The first two, PickPlace and NutAssembly, are from the RoboSuite benchmark, with NutAssembly featuring less initial randomization than the original task. We introduce a third task, TidyUp, to study long-horizon tasks and test the inverse dynamics model’s generalization to unseen environments. We create four environment variants in this domain, denoted as (A, B, C, D). TidyUp task is designed in environment (D), and we collect human demonstrations for TidyUp in the same environment (D). To gauge the inverse dynamics model’s generalization capability, we train two IDMs: IDM-D, based solely on data from environment (D), and IDM-ABC, trained on data from environments (A, B, C). While IDM-D is our default model for experiments, we use IDM-ABC to evaluate generalization in unseen environments.
@@ IV-A1 Manipulation Tasks
We validate our approach and examine the above questions in simulated and real-world tasks. We perform evaluations in three simulated tasks from the RoboSuite simulator [zhu2020robosuite] (see Fig. 3) and two real-world tasks (see Fig. 6):
@@
  • CleanUp. This domain replicates the real-world experimental setup introduced by Nasiriany et al. [nasiriany2022augmenting], involving two tasks: CleanUp-Bin and CleanUp-Stack. The tasks require the robot to push a popcorn box to a serving area and move butter to a target location. See Fig. 6 for the tasks.
We employ a Franka Emika Panda robot with operational space control for all tasks. In simulation and the real world, we opt for 5-DoF control at 20 Hz: 3 for end-effector position $(\delta_{x},\delta_{y},\delta_{z})$, 1 for yaw orientation, and 1 for the gripper. The agent observes wrist-camera and third-person images and receives proprioceptive data. Vision-based policies are trained for all tasks, with only 30 human demonstrations per task. For the training data of IDM, we collect 1M transitions for PickPlace, 3M transitions for NutAssembly, and 5M transitions for TidyUp. Notably, all tasks are completed based on the same set of primitives.
@@ IV-A3 Baselines
[…]
IV-B Experimental Results
TABLE I: Success rates in simulation tasks for ablation studies.
Task Ours No Pretraining Greedy Algo
PickPlace ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{0.967\pm 0.004}}$ $0.463\pm 0.004$ $0.881\pm 0.064$
NutAssembly ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{0.956\pm 0.016}}$ $0.705\pm 0.005$ $0.554\pm 0.080$
TidyUp $\mathbf{0.989\pm 0.005}$ $0.944\pm 0.003$ $0.859\pm 0.004$
@@ IV-B Experimental Results
IV-B1 Quantitative Results
Fig. 4 demonstrates our method’s substantial superiority over all baselines, achieving success rates exceeding 95% across all tasks with remarkable robustness. This showcases the effectiveness of our approach in achieving data-efficient imitation learning through the decomposition of raw sensory demonstrations into concise primitive sequences. Furthermore, our policy often attempts the same primitive type and slightly adjusts primitive parameters after a failure. The results indicate that our primitive policy is more capable of learning this recovery attempt following a failure than baseline policies which require predicting a sequence of actions for recovery. Failures in our method are typically caused by prediction errors in the parameter policy.
PRIME-Learned-Skill outperforms all other baselines, confirming that learned skills are compatible with our PRIME framework. However, the fact that PRIME-Learned-Skill performs worse than PRIME with hard-coded primitives suggests that primitives pretrained from a small set of annotated human demonstrations are suboptimal, leading to a suboptimal low-level controller.
IV-B2 Ablation Studies
We perform ablation studies in all three simulation tasks to study the importance of design choices in our framework. We compare our methods with ablations: 1) without policy pretraining (No Pretraining), 2) utilizing a greedy algorithm to interpret demonstrations into primitive sequences instead of dynamic programming (Greedy Algorithm), where the greedy algorithm selects next primitive by examining all states in the demonstration that come after the current state and choosing the one with the highest probability. As shown in Table I, omitting policy pre-training or substituting dynamic programming with the greedy algorithm leads to decreased performance, highlighting their essential roles.
IV-B3 Model Analysis
@@
IV-C Real-World Evaluation
@@
We evaluate the performance of PRIME against an imitation learning baseline (BC-RNN) on two real-world CleanUp task variants: CleanUp-Bin and CleanUp-Stack. To ensure safe data collection, we perform self-supervised data collection in simulation to train an IDM and apply it to real-world demonstrations. The real-world observations include camera images, object poses, and robot proprioceptive states. Our state-based IDM uses object poses and proprioceptive states, transferring directly to real-world demonstrations. A single IDM segments demonstrations for two tasks within the same domain. Additionally, we develop a visual-based policy to predict motor actions from camera images and robot states. For object poses on the real robot, we use a pose estimator [tremblay2018dope].
The results in Table IV show our method’s notable advantage over BC-RNN, primarily due to its effective task scaffolding. Similar to the results in the simulation, our method also exhibits robust recovery behavior on real robots. Our method has a success rate below 70% in CleanUp-Stack, mainly because stacking demands greater precision and sim2real training for IDM leads to errors.
@@ V Conclusion
We present PRIME, a data-efficient imitation learning approach that decomposes task demonstrations into sequences of primitives and leverages imitation learning to acquire the high-level control policy for sequencing parameterized primitives. While we have already evaluated PRIME with pretrained primitives from annotated demonstrations, a promising direction for future research is to learn a scalable library of low-level skills and compose these diverse skills. These skills can include hard-coded primitives, learned primitives, and skills pretrained from large datasets. This approach holds the potential to facilitate curriculum learning, enabling the progressive acquisition of increasingly complex tasks. A limitation of this study is the use of sim2real experiments for IDM training, which may not be fully applicable to challenging real-world tasks. Extending IDM training to real-world settings is left for future research. Another limitation is that all tasks in this work can be fully decomposed into primitives from the primitive library. Extending our work to include tasks that are not fully decomposable would enhance the generalizability of our framework.
@@ ACKNOWLEDGMENT
The authors would like to thank Yifeng Zhu, Jake Grigsby, Mingyo Seo, Rutav Shah, and Zhenyu Jiang for their valuable feedback. Tian Gao’s visit to UT Austin was supported by IIIS, Tsinghua University. Quantao Yang’s visit to UT Austin was supported by his advisor, Todor Stoyanov, and the Wallenberg AI, Autonomous Systems, and Software Program (WASP). This work has been partially supported by the National Science Foundation (EFRI-2318065, FRR-2145283), the Office of Naval Research (N00014-22-1-2204), UT Good Systems, and the Machine Learning Laboratory.
[…]